# GEO-Bench-2: From Performance to Capability, Rethinking Evaluation in Geospatial AI

### v1

@@ 1 Introduction
Benchmarking is critical in EO, where models must generalize across spatial, temporal, spectral, and sensor domains. While existing benchmarks like PANGAEA [pangaea], Copernicus-Bench [copernicus_bench], and others [fomo, li2025reobench, phileo] have made valuable contributions, they remain limited by narrow modality focus, restrictive licensing, incomplete task coverage, or limited guidelines for reproducible research. GEO-Bench [geobenchv1] introduced structured evaluation protocols but had key gaps: limited task coverage (no detection or instance segmentation), lack of integration with fine-tuning tools, and inability to assess specific model capabilities beyond coarse aggregation.
We present GEO-Bench-2, a comprehensive framework for evaluating GeoFMs across real-world EO applications. Our contributions include: (1) 19 curated datasets with permissive licenses organized into overlapping capability-specific subsets (pixel-wise, detection, multi-temporal, multi-spectral, etc.); (2) geographically balanced splits for scalable, representative evaluation; (3) a prescriptive, but flexible, protocol that ensures fair comparison while allowing for innovation in model finetuning and adaptation; (4) full integration with TerraTorch [terratorch], lowering barriers to entry; and (5) an interactive leaderboard on Hugging Face for community engagement. GEO-Bench-2 enables targeted capability assessment and reproducible comparison, accelerating progress toward general-purpose geospatial intelligence.
@@ 2.1 Dataset Selection and Transformation
There are hundreds of EO labeled datasets [dataset_review]. Through an in-depth investigation, we identified 35 datasets with the potential to meet the criteria below. After experimentation and thorough analysis, we selected 19 that we believe enable high-quality GeoFM evaluation (see Table 2 and full details in Appendix 6.1). Our selection criteria were:
Challenging and Discriminative Datasets must clearly distinguish strong GeoFMs from baseline models. We experimentally validated discriminative power across multiple models, discarding datasets with overlapping performance profiles. For example, most well-trained models can obtain above 98% accuracy on Eurosat [eurosat] or saturate in multi-modal datasets like Sen1Floods11 [sen1floods11].
@@
If needed, datasets were sub-sampled to reduce computational costs. Datasets were formatted following the TACO convention [aybar2025missing], ensuring FAIR compliance [wilkinson2016fair] with self-contained, ML-ready samples. If sub-sampling was applied, a maximum of 20,000 was selected for classification and 4,000 for pixel-wise and detection tasks.
@@
Classification BEN V2 [reben] Land cover S1+S2 10m 20000/4000/4000 19 CDLA-Permissive-1.0
TreeSatAI [treesatai] Tree Species S2 TS 10m 20000/4000/4000 13 $C$-BY-4.0
So2sat [so2sat, geobenchv1] Climate Zones S2 10m 19992/986/986 17 $C$-BY-4.0
Forestnet [forestnet, geobenchv1] Tree Species L8 15m 6464/989/989 12 $C$-BY-4.0
Pixel Regression BioMassters [biomassters] Biomass estim. S1+S2 TS 10m 4000/1000/2000 $\mathbb{R}_{+}$ $C$-BY-4.0
Semantic Segmentation CaFFe [caffe] Glacier zones S1 SAR 10m 4000/1000/2000 4 $C$-BY-4.0
CloudSen12 [cloudsen12+] Cloud/shadow S1+S2 10m 4000/1000/2000 4 CC0
NASA Burn Scars [prithvi2] Burn scars HLS 30m 524/160/120 2 $C$-BY-4.0
Dynamic EarthNet [dynamicearthnet] LULC (temporal) Planet 3m 4000/1000/2000 7 $C$-BY-4.0
FLAIR 2 [flair2] LULC Aerial RGBN+DEM 0.2m 4000/1000/2000 13 Open License 2.0
FTW [fow] Field boundaries S2 10m 4000/1000/2000 2 $C$-BY-SA
KuroSiwo [kurosiwo] Flood extent S1+DEM+Slope 10m 4000/1000/2000 4 MIT
PASTIS (R) [pastisr] Crop type map. S1+S2 10m 1455/482/496 19 $C$-BY-4.0
SpaceNet2 [spacenet] Building Worldview 0.3m 4000/1000/2000 2 $C$-BY-SA-4.0
SpaceNet7 [spacenet] Building Planet 3m 3888/652/1152 2 $C$-BY-SA-4.0
Object Detection EverWatch [everwatch] Bird species Aerial RGB 0.1m 4429/500/196 9 CC0
m-nzcattle [nzcattle] Cattle Aerial RGB 0.1m 524/66/65 2 $C$-BY-4.0
Instance Segmentation PASTIS (R) panoptic [pastisr] Crop type map. S1+S2 10m 1455/482/496 19 $C$-BY-4.0
Substations [substations] Substations S2 10m 4000/500/500 2 $C$-BY-4.0
@@ 3.1 Hyperparameter Optimization
Baseline protocol: Detailed configuration of our experiments is provided in Appendix 6.2. At a glance, we use Optuna [optuna] in TerraTorch Iterate to automate the process, and we encourage users to do so when possible. The optimization is conducted over a fixed hyperparameter space across all models and all datasets, tuning Learning rate and Batch size while keeping other hyperparameters constant. Each HPO and repeated experiment iteration lasted for 50 epochs.
@@ 3.3 Base Model Adaptation
Pixel-wise Tasks: (Segmentation and Regression) We used a UNet decoder for both semantic segmentation and regression tasks, where we fed equally spaced features from the encoder’s output into the UNet [marti2025fine]. For transformer-based models, LearnedFeatureInterpolation, available in TerraTorch, is applied to hierarchically structure the encoder output before passing it to the UNet [unet2015]. This is a technique demonstrated to work well for ViTs compared to other choices (i.e., UPerNet) [marti2025fine].
Object Detection and Instance Segmentation: We used Faster R-CNN [girshick2015fast] and Mask R-CNN [he2017mask], respectively. For both tasks, a FeaturePyramidNetwork [lin2017feature] was applied, following the initial transformation of ViT features as described for pixel-wise tasks.
@@ 3.4 Multi-Spectral Bands, SAR and Multi-Modal Datasets
Multi-Modal Datasets: It remains unclear how much S1 helps when S2 is available [pangaea]. To investigate this, we ran an ablation (Appendix 6.3.6) on three multi-modal models (Terramind, DOFA, Clay), comparing S2-only performance to S1+S2 for multi-modal datasets. For the main results, we report the best outcome per model–dataset pair, which led to using both modalities only for Terramind on BEN V2 and BioMassters (Appendix 6.3.6); all other models use S2 only.
@@ 3.7 Aggregated performance
To obtain a valid aggregation and account for uncertainty introduced by random seeds, we follow the methodology introduced in [geobenchv1]. It is recommended to use the leaderboard or the provided code for computing these scores based on raw results to minimize the chance of mistakes.
Renormalization: To report scores for a capability, we aggregate results across multiple datasets. Simple averaging can distort comparisons, as strong performance on one dataset may overshadow improvements elsewhere [demvsar2006statistical]. Therefore, we normalize each dataset’s scores to [0,1] using a linear transformation with the worst and best result for a given dataset as the reference points [^4] Similar to z-score normalization, but [0,1] improves interpretability.. These normalization factors, computed from our fixed set of models, are published in our Anonymous-Bench-Repo along with tools for calculating final aggregated scores. All users should reuse these factors for consistency.
@@
IQM aggregation: On each normalized bootstrapped iteration, results from each dataset in a capability are averaged across using the interquartile mean (IQM) to discard outliers and obtain a more statistically stable mean [agarwal2021deep].
@@
ConvNext-Large-ImageNet [rw2019timm] ConvNext 230M Supervised ImageNet-22k NA 14M 1 Apache 2.0
ConvNext-XLarge-ImageNet [rw2019timm] ConvNext 390M Supervised ImageNet-22k NA 14M 1 Apache 2.0
DINOv3-ViT-L-SAT [dinov3] ViT 300M Distillation Maxar RGB 0.6 m 493M 1 DINO V3
DINOv3-ConvNext-Large-WEB [dinov3] ConvNext 230M Distillation LVD-1689M NA 1689M 1 DINO V3
Resnet50-DeCUR [wang2024decur] ResNet-50 25M Contrastive Sentinel-2 10 m 1M 1 Apache 2.0
DOFA-ViT-300M [xiong2024dofa] ViT 300M MAE Sentinel-1 and -2, EnMap, Gaofen, Landsat 1-30 m 8M 1 $C$-BY-4.0
Clay-V1 ViT-B [ClayFoundationModel] ViT 86M MAE Landsat 8 and 9, Sentinel-1 and -2, NAIP, LINZ, MODIS 1-30 m 70M 1 Apache 2.0
Satlas-SwinB-Sentinel2 [bastani2023satlaspretrain] Swin 88M Supervised Sentinel-2 10 m NA 1 ODC-BY
Satlas-NAIP [bastani2023satlaspretrain] Swin 88M Supervised NAIP 1 m NA 1 ODC-BY
Prithvi-EO-V2-300M-TL [prithvi2] ViT 300M MAE HLS 30 m 4.2M 4 Apache 2.0
Prithvi-EO-V2-600M-TL [prithvi2] ViT 600M MAE HLS 30 m 4.2M 4 Apache 2.0
TerraMind-V1-Base [jakubik2025terramind] ViT 86M Correlation Sentinel-1 and -2, LULC, DEM, NDVI 10 m 9M 1 Apache 2.0
TerraMind-V1-Large [jakubik2025terramind] ViT 300M Correlation Sentinel-1 and -2, LULC, DEM, NDVI 10 m 9M 1 Apache 2.0
@@ 4.1 Models used in our Experiments
To explore the value of our proposed benchmark, we have chosen models that reflect a diverse range of paradigms, model sizes, and pretraining datasets. They also include foundational general-purpose computer vision models like the latest DINOv3 [dinov3] to explore their generalizability to EO tasks. The characteristics of the benchmarked models are detailed in Table 4. All experiments in the main text are conducted with end-to-end finetuning, meaning the pretrained backbone was unfrozen during the training process.
@@ 4.6 Ablations
Frozen versus Fine-tuned Backbones. We first investigated whether freezing the backbone during training would cause performance degradation and ranking changes relative to full fine-tuning. Contrary to findings in [pangaea], we observed that full fine-tuning consistently outperformed the frozen backbone approach, though the magnitude of performance drop varied across models. This resulted in ranking changes for over 20% of model pairs, as shown in Figure 4.
@@ 5 Discussion and Limitations
Foundation models for EO are advancing rapidly [zhu2024foundations], with a strong appeal for domain-specific applications. However, the field has yet to experience a paradigm shift comparable to that of generative modeling in text or computer vision. While our analysis confirms that larger models tend to perform better, EO has not exhibited the scaling laws observed in LLMs or image generation [hoffmann2022training, li2024scalability].
The LLM community benefits from established benchmarks and centralized infrastructure for tracking progress [white2024livebench]. With GEO-Bench-2, we provide such a framework for the EO domain. A key advantage of the GEO-Bench-2 leaderboard is enabling users to rank models across various capabilities and settings, allowing the community to identify models suited to specific use cases and resource constraints. Our prescriptive protocol is centered on computational budget, repeated experiments, and generalizability, providing consistency while allowing flexibility in how the community adapts foundation models for different capabilities.
Recent work has shown that models pretrained on natural images remain competitive on EO tasks [corley2024revisiting], and well-tuned supervised models can match specialized FMs [xuspecialized]. PANGAEA similarly found that most GeoFMs struggled to beat simpler models under frozen encoder settings [pangaea], with TerraMind being the first exception [jakubik2025terramind].

### v2

@@ 1 Introduction
Benchmarking is critical in EO, where models must generalize across spatial, temporal, spectral, and sensor domains. While existing benchmarks like PANGAEA [42], Copernicus-Bench [59], and others [11, 40, 19] have made valuable contributions, they remain limited by narrow modality focus, restrictive licensing, incomplete task coverage, or limited guidelines for reproducible research. GEO-Bench [38] introduced structured evaluation protocols but had key gaps: limited task coverage (no detection or instance segmentation), lack of integration with fine-tuning tools, and inability to assess specific model capabilities beyond coarse aggregation.
We present GEO-Bench-2, a comprehensive framework for evaluating GeoFMs across real-world EO applications. Our contributions include: (1) 19 curated datasets with permissive licenses organized into overlapping capability-specific subsets (pixel-wise, detection, multi-temporal, multi-spectral, etc.); (2) geographically balanced splits for scalable, representative evaluation; (3) a prescriptive, but flexible, protocol that ensures fair comparison while allowing for innovation in model finetuning and adaptation; (4) full integration with TerraTorch [27], lowering barriers to entry; and (5) an interactive leaderboard on Hugging Face for community engagement. GEO-Bench-2 enables targeted capability assessment and reproducible comparison, accelerating progress toward general-purpose geospatial intelligence.
@@ 2.1 Dataset Selection and Transformation
There are hundreds of EO labeled datasets [49]. Through an in-depth investigation, we identified 35 datasets with the potential to meet the criteria below. After experimentation and thorough analysis, we selected 19 that we believe enable high-quality GeoFM evaluation (see Table 2 and full details in Appendix 6.1). Our selection criteria were:
Challenging and Discriminative Datasets must clearly distinguish strong GeoFMs from baseline models. We experimentally validated discriminative power across multiple models, discarding datasets with overlapping performance profiles. For example, most well-trained models can obtain above 98% accuracy on Eurosat [31] or saturate in multi-modal datasets like Sen1Floods11 [10].
@@
If needed, datasets were sub-sampled to reduce computational costs. Datasets were formatted following the TACO convention [8], ensuring FAIR compliance [62] with self-contained, ML-ready samples. If sub-sampling was applied, a maximum of 20,000 was selected for classification and 5,000 for pixel-wise and detection tasks, with the exception of the Biomasters dataset.
@@
Classification BEN V2 [15] Land cover S1+S2 10m 20000/4000/4000 19 CDLA-Permissive-1.0
TreeSatAI [4] Tree Species S2 TS 10m 20000/4000/4000 13 $C$-BY-4.0
So2sat [65, 38] Climate Zones S2 10m 19992/986/986 17 $C$-BY-4.0
Forestnet [33, 38] Tree Species L8 15m 6464/989/989 12 $C$-BY-4.0
Pixel Regression BioMassters [44] Biomass estim. S1+S2 TS 10m 4011/1739/2776 $\mathbb{R}_{+}$ $C$-BY-4.0
Semantic Segmentation CaFFe [28] Glacier zones S1 SAR 10m 4000/1000/2000 4 $C$-BY-4.0
CloudSen12 [7] Cloud/shadow S1+S2 10m 4000/535/975 4 CC0
NASA Burn Scars [52] Burn scars HLS 30m 524/160/120 2 $C$-BY-4.0
Dynamic EarthNet [53] LULC (temporal) Planet 3m 700/100/200 7 $C$-BY-4.0
FLAIR 2 [21] LULC Aerial RGBN+DEM 0.2m 4049/1022/3022 13 Open License 2.0
FTW [36] Field boundaries S2 10m 4000/1000/2000 2 $C$-BY-SA
KuroSiwo [13] Flood extent S1+DEM+Slope 10m 4000/1000/2000 4 MIT
PASTIS (R) [48] Crop type map. S1+S2 10m 1455/482/496 19 $C$-BY-4.0
SpaceNet2 [57] Building Worldview 0.3m 5186/1461/2961 2 $C$-BY-SA-4.0
SpaceNet7 [57] Building Planet 3m 3888/652/1152 2 $C$-BY-SA-4.0
Object Detection EverWatch [24] Bird species Aerial RGB 0.1m 4429/500/196 9 CC0
m-nzcattle [1] Cattle Aerial RGB 0.1m 524/66/65 2 $C$-BY-4.0
Instance Segmentation PASTIS (R) panoptic [48] Crop type map. S1+S2 10m 1455/482/496 19 $C$-BY-4.0
Substations [35] Substations S2 10m 4000/500/500 2 $C$-BY-4.0
@@ 3.1 Hyperparameter Optimization
Baseline protocol: Detailed configuration of our experiments is provided in Appendix 6.2. At a glance, we use Optuna [5] in TerraTorch Iterate to automate the process, and we encourage users to do so when possible. The optimization is conducted over a fixed hyperparameter space across all models and all datasets, tuning Learning rate and Batch size while keeping other hyperparameters constant. Each HPO and repeated experiment iteration lasted for 50 epochs.
@@ 3.3 Base Model Adaptation
Pixel-wise Tasks: (Segmentation and Regression) We used a UNet decoder for both semantic segmentation and regression tasks, where we fed equally spaced features from the encoder’s output into the UNet [43]. For transformer-based models, LearnedFeatureInterpolation, available in TerraTorch, is applied to hierarchically structure the encoder output before passing it to the UNet [47]. This is a technique demonstrated to work well for ViTs compared to other choices (i.e., UPerNet) [43].
Object Detection and Instance Segmentation: We used Faster R-CNN [26] and Mask R-CNN [30], respectively. For both tasks, a FeaturePyramidNetwork [41] was applied, following the initial transformation of ViT features as described for pixel-wise tasks.
@@ 3.4 Multi-Spectral Bands, SAR and Multi-Modal Datasets
Multi-Modal Datasets: It remains unclear how much S1 helps when S2 is available [42]. To investigate this, we ran an ablation (Appendix 6.3.6) on three multi-modal models (Terramind, DOFA, Clay), comparing S2-only performance to S1+S2 for multi-modal datasets. For the main results, we report the best outcome per model–dataset pair, which led to using both modalities only for Terramind on BEN V2 and BioMassters (Appendix 6.3.6); all other models use S2 only.
@@ 3.7 Aggregated performance
To obtain a valid aggregation and account for uncertainty introduced by random seeds, we follow the methodology introduced in [38]. It is recommended to use the leaderboard or the provided code for computing these scores based on raw results to minimize the chance of mistakes.
Renormalization: To report scores for a capability, we aggregate results across multiple datasets. Simple averaging can distort comparisons, as strong performance on one dataset may overshadow improvements elsewhere [18]. Therefore, we normalize each dataset’s scores to [0,1] using a linear transformation with the worst and best result for a given dataset as the reference points [^4] Similar to z-score normalization, but [0,1] improves interpretability.. These normalization factors, computed from our fixed set of models, are published in our Anonymous-Bench-Repo along with tools for calculating final aggregated scores. All users should reuse these factors for consistency.
@@
IQM aggregation: On each normalized bootstrapped iteration, results from each dataset in a capability are averaged across using the interquartile mean (IQM) to discard outliers and obtain a more statistically stable mean [2].
@@
ConvNext-Large-ImageNet [61] ConvNext 230M Supervised ImageNet-22k NA 14M 1 Apache 2.0
ConvNext-XLarge-ImageNet [61] ConvNext 390M Supervised ImageNet-22k NA 14M 1 Apache 2.0
DINOv3-ViT-L-SAT [50] ViT 300M Distillation Maxar RGB 0.6 m 493M 1 DINO V3
DINOv3-ConvNext-Large-WEB [50] ConvNext 230M Distillation LVD-1689M NA 1689M 1 DINO V3
Resnet50-DeCUR [58] ResNet-50 25M Contrastive Sentinel-2 10 m 1M 1 Apache 2.0
DOFA-ViT-300M [63] ViT 300M MAE Sentinel-1 and -2, EnMap, Gaofen, Landsat 1-30 m 8M 1 $C$-BY-4.0
Clay-V1 ViT-B [16] ViT 86M MAE Landsat 8 and 9, Sentinel-1 and -2, NAIP, LINZ, MODIS 1-30 m 70M 1 Apache 2.0
Satlas-SwinB-Sentinel2 [9] Swin 88M Supervised Sentinel-2 10 m NA 1 ODC-BY
Satlas-NAIP [9] Swin 88M Supervised NAIP 1 m NA 1 ODC-BY
Prithvi-EO-V2-300M-TL [52] ViT 300M MAE HLS 30 m 4.2M 4 Apache 2.0
Prithvi-EO-V2-600M-TL [52] ViT 600M MAE HLS 30 m 4.2M 4 Apache 2.0
TerraMind-V1-Base [34] ViT 86M Correlation Sentinel-1 and -2, LULC, DEM, NDVI 10 m 9M 1 Apache 2.0
TerraMind-V1-Large [34] ViT 300M Correlation Sentinel-1 and -2, LULC, DEM, NDVI 10 m 9M 1 Apache 2.0
@@ 4.1 Models used in our Experiments
To explore the value of our proposed benchmark, we have chosen models that reflect a diverse range of paradigms, model sizes, and pretraining datasets. They also include foundational general-purpose computer vision models like the latest DINOv3 [50] to explore their generalizability to EO tasks. The characteristics of the benchmarked models are detailed in Table 4. All experiments in the main text are conducted with end-to-end finetuning, meaning the pretrained backbone was unfrozen during the training process.
@@ 4.6 Ablations
Frozen versus Fine-tuned Backbones. We first investigated whether freezing the backbone during training would cause performance degradation and ranking changes relative to full fine-tuning. Contrary to findings in [42], we observed that full fine-tuning consistently outperformed the frozen backbone approach, though the magnitude of performance drop varied across models. This resulted in ranking changes for over 20% of model pairs, as shown in Figure 4.
@@ 5 Discussion and Limitations
Foundation models for EO are advancing rapidly [66], with a strong appeal for domain-specific applications. However, the field has yet to experience a paradigm shift comparable to that of generative modeling in text or computer vision. While our analysis confirms that larger models tend to perform better, EO has not exhibited the scaling laws observed in LLMs or image generation [32, 39].
The LLM community benefits from established benchmarks and centralized infrastructure for tracking progress [60]. With GEO-Bench-2, we provide such a framework for the EO domain. A key advantage of the GEO-Bench-2 leaderboard is enabling users to rank models across various capabilities and settings, allowing the community to identify models suited to specific use cases and resource constraints. Our prescriptive protocol is centered on computational budget, repeated experiments, and generalizability, providing consistency while allowing flexibility in how the community adapts foundation models for different capabilities.
Recent work has shown that models pretrained on natural images remain competitive on EO tasks [17], and well-tuned supervised models can match specialized FMs [64]. PANGAEA similarly found that most GeoFMs struggled to beat simpler models under frozen encoder settings [42], with TerraMind being the first exception [34].